OpenStudy (anonymous):

What is the sum of a 30–term arithmetic sequence where the first term is 72 and the last term is –102?

OpenStudy (anonymous):

The sum is \(n\frac{a_1+a_n}{2}=30\frac{72-102}{2}=-450.\)